# MixLLM: LLM Quantization with Global Mixed-precision between Output-features and Highly-efficient System Design

### v1

@@ 3.2 Global Precision Search Algorithm
As discussed in Sec.3.1, MixLLM determines the precision of all output features in all layers globally rather than locally. It identifies the salience of these features with respect to the final loss of the model, and assigns larger bit-width to the features leading to larger loss.
Specifically, it calculates the salience $S$ of a channel $c$ as:
@@
which is the distance of the model’s loss between quantizing and not quantizing this single channel. In Eq.1, $l()$ is the loss function of the model w.r.t. a single channel, $c_{q}$ is the quantized weight of the channel and $c_{0}$ is the original weight. Note that it regards other neurons except $c$ as constant in $l()$.
We use the Taylor Expansion method to estimate the loss function $l(c)$ (similar with the existing quantization works, ignoring the high-order items):
@@
where $g=\mathbb{E}[\frac{\partial}{\partial c}l(c)]$ is the loss’s gradient w.r.t. the channel, and $H=\mathbb{E}[\frac{\partial^{2}}{\partial c^{2}}l(c)]$ is the second-order gradient (i.e., Hessian matrix) w.r.t. the channel.
It is infeasible to calculate the Hessian matrix as it is too costly. We approximate the Hessian $H$ with the (empirical) Fisher information matrix $\mathit{F}$ on the calibration dataset $\mathit{D}$:
$H\approx\mathit{F}=\frac{1}{\lvert\mathit{D}\rvert}\sum_{d\in\mathit{D}}{g_{d}g_{d}^{T}}$ (3)
Note that $\mathit{F}$ is w.r.t. a channel, differing from the diagonal Fisher information matrix in the recent works that ignores any cross-neuron interactions [23, 21].
@@
We do not ignore the first order information during the calculation, differing from OBD [24] and many recent quantization works [14, 11, 21]. This is because the first order factor can be more significant in the estimation in Eq.4, as the estimated second order factor is the square of the first order factor divided by two for each sample. Considering that $g$ can be very small for the well pretrained models and the delta of the quantized weight is usually not large, the first order factor can be larger than the second order one. Besides, what we require is the loss itself rather than the arguments of the loss function, and thus we do not need to ignore the first order factor to simplify the arguments calculation.
@@
1:Linear layer number $L$, weight and gradient of all linear layers ($W_{i}\in\mathbb{R}^{O,I}$, $G_{i}\in\mathbb{R}^{O,I}$ for layer $i$).
2:Global channel index with large and small bit width (largebit_channels, smallbit_channels).
3:$S_{global}$leftarrow []
4:for $i=1,2,...,L$ do
5:     $W_{i}^{delta}$leftarrow quantize($W_{i}$) - $W_{i}$
6:     $S_{1st}$leftarrow sum($G_{i}\odot W_{i}^{delta}$, dim=1) $\triangleright$ Per-channel dot product between $G_{i}$ and $W_{i}^{delta}$
7:     $S_{2nd}\leftarrow 0.5*(S_{1st})^{2}$
8:     $S\leftarrow\lvert S_{1st}+S_{2nd}\rvert$triangleright $S\in\mathbb{R}^{O}$, the salience of the $O$ channels
9:     for $channel\_id=1,2,..,O$ do $\triangleright$ Log the salience of each output channel of this layer
10:         $S_{global}$.append(tuple($i$, $channel\_id$, $S[channel\_id]$))
11:sort($S_{global}$) $\triangleright$ Sort according to the salience, descending
12:largebit_channels, smallbit_channels $\leftarrow$ $S_{global}[:N_{largebit}]$, $S_{global}[N_{largebit}:]$
Algo.1 illustrates the procedure of the global precision search. It calculates the salience of all the output channels of all linear layers and sort them in descending order globally. Given the global threshold $N_{largebit}$ as the number of large-bit precision channels, the first $N_{largebit}$ channels are intended to be quantized with 8-bit, and the other channels will be quantized with smaller bit-width (i.e., 4-bit in this paper). Any quantization methodologies (e.g., GPTQ, clip search) can be applied independently to these two disjoint parts of channels. Note that we calculate the salience of the channels in one pass rather than iterative identifying the high-salience parts in a smaller step, as we observe the single-pass method show similar results with the iterative method and saves a lot of computation overhead than the latter.
@@ 3.3 Efficient Quantization Computation System
Two-step dequantization to make use of int8 Tensor Core. As for the W4A8 computation, the dequantized weight and activations are $(W_{q}-z)s_{w}$ and $A_{q}s_{a}$, where $W_{q}$ and $z$ are uint4 datatype, $A_{q}$ is int8 datatype, and $s_{w}$ and $s_{a}$ are float16 datatype. Directly dequantizing the tensors into float16 datatype before the MatMul computation will prevent us using the fast 8-bit Tensor Core on the GPU. Instead, MixLLM uses a two step dequantization within each group. Specifically, MixLLM first partially dequantizes the weight into $(W_{q}-z)$, and then multiply it by $A_{q}$ with the 8-bit Tensor Core. Finally, it multiplies this MatMul result by the two scales within each group. Note that we use int8 datatype for $(W_{q}-z)$ so that there is no overflow problem.
@@
Fast Int to Float conversion with partially fusing into Tensor Core instruction. In the above two-step dequantization computation, the step 2 is the MatMul between the integer tensor $A_{q}(W_{q}-z)$ and the float tensor $s_{a}s_{w}$. It requires the integer to float conversion (I2F) before the multiply operation. The I2F instruction is expensive on the modern GPUs. Instead, we make use of the range-dependent fast I2F transformation to convert the I2F instruction into two add/sub instructions [^4] CUTLASS [9] also has an implementation of fast I2F for general purpose.. Specifically, it is based on the fact that there exists a certain range where an integer value’s binary is the same as a corresponding float binary. As shown in Fig.3, the binary with the first 9 bits as 010010110 represents a series of consecutive int32 and float32 values, respectively. We can add a bias to an integer value to make it within this consecutive range, and then subtract a corresponding bias in float (same underling binary) to restore its value in float type. We take the middle value in this range as the bias to maximize the data range that can be safely converted, whose hexadecimal number is 0x4b400000 (i.e., in the remaining 23 bits, the first bit is 1 and the other bits are 0). This allows to convert a consecutive range of $2^{23}$ int32 numbers to float32. The range of dot product of $k$ int8 elements is $2^{16}k$, thus the above fast I2F conversion allows the $k$ value up to 128. We use quantization group size as 128 and can use the fast I2F safely:
1// bias_int = as_int(0x4b400000), bias_fp = as_float(0x4b400000);
@@
We further fuse the integer subtraction into the Tensor Core mma (Matrix Multiply-Accumulate) instruction. The mma instruction computes $D=AB+D$ during the MatMul computation. We initialize the accumulator $D$ as the bias_int before MatMul computation of each quantization group, and will only need to subtract the bias_fp after the MatMul. In another word, the expensive I2F is converted into a single float subtraction. The above I2F simplification brings more than 20 TOPS performance improvement for the 512/4096/4096 (M/N/K) quantized MatMul computation on an A100 GPU.
Figure 4: The GPU kernel software pipeline of group-wise W4A8/W8A8 quantized MatMul. It assumes perfect overlapping. G2S: load global to shared memory; S2R: load shared memory to register; MMA: matrix multiply-accumulation; I2F: integer to float conversion; deq: dequantize; acc: accumulate.
@@
End-to-end software pipeline of the quantized linear kernel on the GPU. Fig.4 shows the software pipeline of the quantized kernel. Beside the basic warp tile and block tile, we introduce the quantization group tile for the fast I2F and per-group scale multiplication. It uses two output buffers for the output accumulation at register level, one for the per-group accumulation, and the other for the global accumulation. This allows to apply the per-group scales on the group-level buffer. We initialize the group buffer with the bias_int at the beginning of the group tile, and subtract bias_fp at the end of the group tile. As for the two-step dequantization, the first step is within the warp tile where each input element will subtract the zero-point before feeding into MMA, the second step is at the end of the group tile by multiplying the scale. We use the vectorized intrinsic to perform four int8 subtract in a single instruction (vsub4[27]. Besides, to improve the global memory loading efficiency, we prepack the memory layout of the weight tensor ahead-of-time to avoid the runtime permutation of the input elements. In general, this software pipeline can overlap the memory loading and computation, and the dequantization computation with SIMT Core and the MatMul computation with Tensor Core to the best, and minimizes the overhead of group-wise dequantization.
@@
Parallel execution of sub-problems of different bit-width. As for the execution shown in Fig.1, MixLLM executes different sub-problems in parallel on the GPU with CUDA Graph. Finally, the MatMul execution of the two parts write to the same target tensor with different channel indices to generate the final output. We implement this function with the fused epilogue of MatMul to scatter the output to the corresponding indices, which is basically costless.
@@ 4 Evaluation
4.1 Setup
As for MixLLM evaluation in this paper, we use 0%, 10%, 20%, 50% and 100% percent of 8-bit based on the 4-bit quantization, respectively. Meanwhile, we use 8-bit for activation quantization. Both the weight and activation are group-wise quantized with group size 128. The 4-bit part is asymmetric quantized and the 8-bit part (including that in weight) is symmetric, which is a good trade-off between accuracy and system efficiency. Note that any other bit-width percentage configuration can be used for real scenarios to trade-off memory usage, system efficiency and accuracy in practice. We enable GPTQ in MixLLM for the all models. We also apply clip search for all the models except for the 32B, 70B and 72B models as the clip search is time consuming for large models.
Baselines and configurations. We compare MixLLM with the state-of-the-art (SOTA) quantization solutions of both weight-only and weight-activation methods. As for the weight only quantization, we compare MixLLM with the basic round-to-nearst (RTN) 4-bit and 5-bit quantization, and the production-level SOTA GPTQ [14] and AWQ[26]. As for the weight-activation quantization, we compare MixLLM with the most widely used SmoothQuant [45] and the recent SOTA QoQ [27] and QuaRot[3] (of both W4A4 and W4A8). The 8-bit tensors are all symmetric quantized in all baselines and MixLLM. We also compare the perplexity with SqueezeLLM[21], OminiQuant[37], AffineQuant[29], Atom[50] and SpinQuant[28] according to their reported numbers.
We make use of AutoGPTQ lib [5] (v0.7.1) to evaluate GPTQ, AutoAWQ lib [4] (v0.2.7) to evaluate AWQ, lmquant lib [33] (commit 9d62c5c) to evaluate SmoothQuant and QoQ, and the official repo to evaluat QuaRot. We enable the reorder trick for GPTQ evaluation, and use asymmetric and group size 128 for both GPTQ and AWQ. We follow the official configurations in lmquant to use 0.85/0.15 as the alpha/beta parameter for SmoothQuant, and 0.3/0.7 for QoQ. We use symmetric and per-channel/token quantization in QuaRot, following the configuration in its paper. We disable the KV quantization of QoQ and QuaRot in our experiments to make the comparison fair.
Models and Datasets. We evaluate MixLLM and the baselines on a variety of widely used LLMs of different sizes, ranging from 0.5B to 72B. The models include Llama 3.1 8B and 70B [32], Llama 3.2 1B, Qwen2.5 0.5B, 1.5B, 7B and 32B [17], and Mistral 7B v0.3 [20].
We use wikitext2 dataset [31] as the calibration set for GPTQ and MixLLM. We use the default pile dataset [34] as the calibration dataset for AWQ, SmoothQuant and QoQ, to enable their better performance. GPTQ, AWQ and MixLLM uses 128 samples with sequence length of 2048 for calibration. SmoothQuant and QoQ uses 64 samples with sequence length of 1024 for calibration (larger dataset results in OOM in our experiment).
Metrics. As for the algorithm accuracy, we compare the perplexity (ppl) between all the baselines on wikitext2 and C4 [35] dataset. Meanwhile, we compare a set of popular downstream tasks on Llama 3.1 8B, Qwen2.5 7B, and Mixtral 7B v0.3, including BBH [39], GPQA [36], MMLU-Pro [41], MuSR [38], ARC challenge [6], and HellaSwag [49]. We use lm-eval [15] for the downstream tasks evaluation, for which the task names are leaderboard_bbh, leaderboard_gpqa, leaderboard_mmlu_pro, leaderboard_musr, arc_challenge, and hellaswag. We use the average number for each of the tasks.
We conduct the system experiments on NVIDIA A100 (80G) GPUs with CUDA 12.1. We use PyTorch 2.4.1 and transformers 4.45.2.
4.2 Perplexity Evaluation
Table 1: Perplexity evaluation ($\downarrow$) on wikitext2/c4 (gray for c4), sequence length 2048. NA means no support. Abn means the value is too large ($>10^{5}$). For MixLLM, pn means $n\%$ 8-bit.
@@ 4.2 Perplexity Evaluation
Tab.1 shows the perplexity on Wikitext2 and C4 dataset for the commonly used open source LLMs, of different baselines. It shows that:
  • Using 4.4 bits of weights with MixLLM can achieve the similar accuracy with 5 bits RTN weight-only quantization, even with 8-bit activation quantization enabled in MixLLM. This is mainly because MixLLM assigns the high-salience output channels with larger bit-widths than the uniform 5-bit solution.
  • As for the weight-only quantization baselines, MixLLM W4.4A8 outperforms the production SOTA solutions GPTQ and AWQ, with only 10% more bit-width, and even with 8-bit activation quantization enabled in MixLLM. Meanwhile, the RTN W5A16 method also outperforms GPTQ and AWQ, which means a slightly larger bit-width can defeat the well tuned smaller bit-width easily. MixLLM W4.4A8 benefits from the larger bits on the top 10% output features with high salience.
  • As for the weight-activation quantization baselines, MixLLM W4.4A8 shows a comparable accuracy with SmoothQuant with much smaller bit-width (60% of that in SmoothQuant). MixLLM W4.4A8 shows better accuracy than QoQ and QuaRot with only 10% larger bit-width. It shows MixLLM achieves a good balance of memory consumption and accuracy.
  • Note that MixLLM W8A8 quantization shows nearly lossless performance compared to the float16 baseline. This is part of the motivation that MixLLM uses group-wise quantization for the activation.
LLaMA 2 FP16
SqueezeLLM
W4A16 0.45%
OminiQuant
W4A16/W4A4
AfineQuant
W4A16/W4A4
Atom
W4A4 128 outliers
SpinQuant
W4A8
MixLLM
W4.4A8
7B 5.47 5.57 5.58/14.26 5.58/12.69 6.03 5.7 5.55
13B 4.88 4.96 4.95/12.30 4.95/11.45 5.27 5.0 4.93
Comparison with More Related Works We compare MixLLM with more recent quantization works according to the reported numbers in their papers (Tab.2), showing that MixLLM achieves superior accuracy to a broad range of related works with similar memory consumption.

### v2

@@ 3.2 Global Precision Search Algorithm
MixLLM determines the precision of all output features in all layers globally. It identifies the salience of these features with respect to the final loss of the model, and assigns larger bit-width to the features leading to larger loss. Specifically, it calculates the salience $S$ of a channel $c$ as:
@@
which is the distance of the model’s loss between quantizing and not quantizing channel $c$. In Eq.1, $l()$ is the loss function of the model w.r.t. a single channel, $c_{q}$ is the quantized weight of the channel and $c_{0}$ is the original one.
We use the Taylor Expansion method to estimate the loss function $l(c)$, ignoring the high-order items:
@@
where $g=\mathbb{E}[\frac{\partial}{\partial c}l(c)]$ is the loss’s gradient w.r.t. the channel, and $H=\mathbb{E}[\frac{\partial^{2}}{\partial c^{2}}l(c)]$ is the second-order gradient (i.e., Hessian matrix). It is infeasible to calculate the Hessian matrix. We approximate the Hessian with the (empirical) Fisher information matrix (FIM) $F$ on the calibration dataset $\mathcal{D}$:
$H\approx F=\frac{1}{\lvert\mathcal{D}\rvert}\sum_{d\in\mathcal{D}}{g_{d}g_{d}^{T}}$ (3)
Note that $F$ is w.r.t. a channel, differing from the diagonal FIM in the recent works that ignores any cross-neuron interactions Kwon et al. (2022); Kim et al. (2024).
@@
We do not ignore the first order information during the calculation, differing from the recent quantization works Frantar et al. (2022); Dettmers et al. (2024); Kim et al. (2024). Note that what we require is the loss itself rather than the arguments of the loss function, thus we do not need to ignore the first order factor to simplify the arguments calculation.
@@
0: Weight and gradient of all linear layers ($W_{i}\in\mathbb{R}^{O\times I}$, $G_{i}\in\mathbb{R}^{O\times I}$ for layer $i\in[1..L]$). Total number of output channels with large bit-width ($T$).
0: Channel index for large and small bit-width ($\mathcal{C}^{lg}$ and $\mathcal{C}^{sm}$).
[…]
2:for $i=1$ to $L$ do
3:  $W_{i}^{delta}$ $\leftarrow$ quantize($W_{i}$) - $W_{i}$
4:  $S^{1st}$ $\leftarrow$ sum($G_{i}\odot W_{i}^{delta}$, dim=1)
5:  $S^{2nd}\leftarrow 0.5*(S^{1st})^{2}$
6:  $S\leftarrow\lvert S^{1st}+S^{2nd}\rvert \triangleright$ $O$ channels’ salience $\in\mathbb{R}^{O}$
7:  for $cid=1$ to $O$ do
8:   $\mathcal{C}^{global}$ $\leftarrow$ $\mathcal{C}^{global}$ $\cup$ ( tuple(ii, cidcid, $S_{cid}$) )
9:  end for
10:end for
11: Sort $\mathcal{C}^{global}$ by salience in descending order.
[…]
Algo.1 illustrates the procedure of the global precision search. It calculates the salience of all the output channels of all linear layers and sort them in descending order. Given the global threshold $T$ as the number of large-bit precision channels, the first $T$ channels are intended to be quantized with 8-bit, and the other channels will be quantized with smaller bit-width (i.e., 4-bit in this paper).
@@ 3.3 Efficient Quantization Computation System
Figure 3: The GPU kernel software pipeline of group-wise W4A8/W8A8 quantized MatMul. It assumes perfect overlapping. G2S: load global to shared memory; S2R: load shared memory to register; MMA: matrix multiply-accumulation; I2F: integer to float conversion; deq: dequantize; acc: accumulate. While this pipeline is modeled on the NVIDIA A100 architecture, its fundamental principles remain applicable to subsequent generations, such as Hopper and Blackwell, subject to minor architectural refinements. For instance, newer architectures can utilize the Tensor Memory Accelerator (TMA) to load activation tensors directly, bypassing registers before they reach the Tensor Cores. Furthermore, these architectures support warp specialization for memory loading as an alternative to a uniform execution scheme.
Two-step dequantization to make use of int8 Tensor Core. As for the W4A8 computation, the dequantized weight and activations are $(W_{q}-z)s_{w}$ and $A_{q}s_{a}$, where $W_{q}$ and $z$ are uint4 datatype (4-bit unsigned integer), $A_{q}$ is int8 datatype, and $s_{w}$ and $s_{a}$ are float16 datatype. Directly dequantizing the tensors into float16 datatype before the MatMul computation will prevent us using the fast 8-bit Tensor Core on the GPU. Instead, MixLLM uses a two step dequantization within each group. Specifically, MixLLM first partially dequantizes the weight into $(W_{q}-z)$, and then multiply it by $A_{q}$ with the 8-bit Tensor Core. Finally, it multiplies this MatMul result by the two scales within each group. Note that we use int8 datatype for $(W_{q}-z)$ so that there is no overflow problem.
@@
Fast I2F with partially fusing into Tensor Core instruction. In the above two-step dequantization computation, the step 2 is the multiplication between integer and float tensor, requiring integer to float conversion (I2F). As I2F instruction is expensive on the modern GPUs, we make use of the range-dependent fast I2F transformation to convert the I2F instruction into two add/sub instructions. Specifically, it is based on the fact that there exists a certain range where an integer value’s binary is the same as a corresponding float binary. As shown in Fig.4, the binary with the first 9 bits as 010010110 represents a series of consecutive int32 and float32 values, respectively. We can add a bias to an integer value to make it within this consecutive range, and then subtract a corresponding bias in float to restore its value in float. We take the middle value in this range as the bias to maximize the data range that can be safely converted, whose hexadecimal number is $mid=\texttt{0x4b400000}$ (i.e., in the remaining 23 bits, the first bit is 1 and the other bits are 0). This allows to convert a consecutive range of $2^{23}$ int32 numbers to float32. The range of dot product of $k$ int8 elements is in $2^{16}k$, thus the above fast I2F conversion allows the $k$ value to be 128. We use quantization group size as 128 and can use the fast I2F safely:
1// bias_int = as_int(mid), bias_fp = as_float(mid);
@@
We further fuse the integer subtraction into the Tensor Core mma (Matrix Multiply-Accumulate $D=AB+D$) instruction. We initialize the accumulator $D$ as the bias_int before MatMul computation of each quantization group, and will only need to subtract the bias_fp after the MatMul. In another word, the expensive I2F is converted into a single float subtraction. The above I2F simplification brings more than 20 TOPS performance improvement for the 512/4096/4096 (M/N/K) quantized MatMul computation on an A100 GPU.
@@
End-to-end software pipeline of the quantized linear kernel on the GPU. Fig.3 shows the software pipeline of the quantized kernel. Besides the basic warp tile and block tile, we introduce the quantization group tile for the per-group quantized computation. It uses two output buffers for the output accumulation at register level, one for the per-group accumulation, and the other for the global accumulation. This allows to apply the per-group scales on the group-level buffer. We initialize the group buffer with the bias_int at the beginning of the group tile, and subtract bias_fp at the end of the group tile. As for the two-step dequantization, the first step is within the warp tile where each input element will subtract the zero-point before feeding into MMA, the second step is at the end of the group tile by multiplying the scale. We use the vectorized intrinsic to perform four int8 subtract in a single instruction (vsub4Lin et al. (2025). Besides, to improve the global memory loading efficiency, we prepack the memory layout of the weight tensor ahead-of-time to avoid the runtime permutation of the input elements. This software pipeline can overlap the memory loading, the dequantization computation with SIMT Core, and the MatMul computation with Tensor Core to the best, and minimizes the overhead of group-wise dequantization.
3.4 Parallel Execution of Sub-problems of Different Bit-width
@@
As for the execution shown in Fig.1, MixLLM executes different sub-problems in parallel on the GPU with CUDA Graph. Finally, the MatMul execution of the two parts write to the same target tensor with different channel indices to generate the final output. We implement this function with the fused epilogue of MatMul to scatter the output to the corresponding indices, which is basically costless.
3.5 Quantization-aware Roofline Discussion
To systematically analyze the performance bounds of MixLLM quantization schemes, we extend the traditional Roofline model to account for the heterogeneous compute pipelines on NVIDIA GPUs. Using the NVIDIA A100 (80GB SXM) as our target architecture, we evaluate the system against its peak hardware limits: 2039 GB/s of High Bandwidth Memory (HBM2e) throughput, 624 TOPS of dense INT8 Tensor Core compute, and 19.5 TFLOPS of FP32 SIMT compute. Introducing sub-byte quantization and asymmetric zero-points fundamentally alters the arithmetic intensity (operations per byte) and introduces multiple competing compute ceilings across the execution units.
Memory Traffic. From a memory bandwidth perspective, arithmetic intensity is governed by the quantized data footprint and associated metadata. For a group size of $G=128$, the W8A8 symmetric scheme fetches 8-bit weights, 8-bit activations, and a 16-bit scale, yielding an effective weight footprint of $1+2/G$ bytes per element. The W4A8 asymmetric scheme compresses weights to 4 bits, requiring a 16-bit scale and a 4-bit zero-point, yielding an effective footprint of $0.5+2.5/G$ bytes per weight. Consequently, W4A8 drastically shifts the memory-bound slope upward, allowing significantly higher attainable throughput in bandwidth-constrained regimes, such as small batch decoding. This motivates quantization with lower bits for weight in MixLLM.
Compute Ceilings: Tensor Core vs. SIMT Dequantization. In the compute-bound regime, performance is dictated by the interplay between the Tensor Cores (handling the GEMM) and the SIMT cores (handling dequantization).
For both W8A8 and W4A8, output scaling requires a transition from the Tensor Core’s int32 accumulators to the SIMT core’s float32 pipeline. For every group size of $G=128$, the Tensor Cores perform 128 INT8 multiply-accumulate (MAC) operations (256 operations total) to produce a single int32 partial sum. Before the FP32 scales ($s_{w}$ and $s_{a}$) can be applied, the SIMT cores must first issue an int32-to-float32 cast instruction. Therefore, the post-GEMM workload per group consists of this explicit cast instruction followed by the FP32 scale multiplications. Even with this added casting overhead, the application’s required operational ratio—256 INT8 operations per 3 SIMT instructions (cast + multiplies)—remains well below the A100 hardware’s provisioned ratio of 32:1 (624 INT8 TOPS vs. 19.5 FP32 TFLOPS), thanks to the fast I2F and partial fusion methodology in MixLLM. Because the SIMT requirement is so light, this post-GEMM casting and scaling can be hidden behind the Tensor Core accumulation. This further motivates the group-wise 8-bit quantization than per-channel/per-token solutions.
However, W4A8 introduces a heavier pre-GEMM dequantization step:
$Y=((W_{q}-z)\times A_{q})\cdot s_{w}\cdot s_{a}$
Before the Tensor Cores can execute the INT8 MACs, the 4-bit weights must be unpacked and the zero-point subtracted. By leveraging the vsub4 intrinsic, the hardware can perform 4 int8 subtractions within a single SIMT instruction, highly optimizing the $(W_{q}-z)$ computation. Despite this instruction-level parallelism, the SIMT integer pipeline must still continuously feed the Tensor Cores. Because this pre-GEMM workload scales at $O(N^{2})$ and competes for instruction fetch and register bandwidth concurrently with the $O(N^{3})$ Tensor Core math, it creates an auxiliary, lower compute ceiling. Thus, the effective peak compute for W4A8 is strictly lower than the absolute 624 TOPS achieved by W8A8.
Intersection Point (Ridge Point) Analysis. The traditional Roofline ridge point defines the intersection between memory bandwidth and peak compute, marking the transition from a memory-bound to a compute-bound workload. For pure INT8 Tensor Core execution on the A100, this intersection occurs at roughly 306 OPs/Byte (624 TOPS / 2039 GB/s). Because the W4A8 scheme is constrained by the SIMT vsub4 dequantization ceiling rather than the peak Tensor Core ceiling, its horizontal Roofline ceiling drops. Consequently, the intersection between the W4A8 memory traffic line and its compute ceiling shifts to the left. This leftward shift indicates that W4A8 workloads saturate their maximum achievable compute at a lower arithmetic intensity than W8A8, entering the compute-bound regime earlier due to SIMT instruction overhead.
@@ 4 Evaluation
Table 1: Perplexity evaluation ($\downarrow$) on wikitext2/c4 (gray for c4), sequence length 2048. NA means no support. Abn means the value is too large ($10^{5}$). For MixLLM, pn means $n\%$ 8-bit.
@@
4.1 Setup
As for MixLLM evaluation, we use 0%, 10%, 20%, 50% and 100% of 8-bit based on the 4-bit quantization, respectively. Meanwhile, we use 8-bit for activation quantization. Both the weight and activation are group-wise quantized with group size 128. The 4-bit part is asymmetric quantized and the 8-bit part (including that in weight) is symmetric. Similar to the recent work Lin et al. (2025); Liu et al. (2024), we enable GPTQ and clipping in MixLLM.
Baselines and configurations. In Sec.4.2 and Sec.4.4, we compare to the pure PTQ methods of weight-activation quantization, neither the QAT nor the methods training the quantization parameters (e.g., training the rotation matrix like in SpinQuantLiu et al. (2024)). Specifically, we compare MixLLM with the SOTA PTQ weight-activation methods, including the most widely used SmoothQuant Xiao et al. (2023) and the recent SOTA QuaRotAshkboos et al. (2024) (of both W4A4 and W4A8) and QServe Lin et al. (2025). The 8-bit tensors are all symmetric quantized in all baselines. We use symmetric and per-channel/token quantization in QuaRot, following the setting in its paper. We follow the official configurations to use 0.85/0.15 as the alpha/beta parameter for SmoothQuant, and 0.3/0.7 for QServe. We disable the KV quantization of QuaRot and QServe in our experiments to make the comparison fair. In Appendix.4.6, We also compare the ppl with GPTQ Frantar et al. (2022), AWQLin et al. (2024), SqueezeLLMKim et al. (2024), OmniQuantShao et al. (2024), AffineQuantMa et al. (2024), AtomZhao et al. (2024) and SpinQuantLiu et al. (2024).
Models and Datasets. We evaluate MixLLM and the baselines on Llama 3.1 8B and 70B Meta (Cited 2024), Llama 3.2 1B, Qwen2.5 0.5B, 1.5B, 7B and 32B Group (Cited 2024), and Mistral 7B v0.3 Jiang et al. (2023). We use wikitext2 Merity et al. (2017) as the calibration set for MixLLM, and the default pile dataset MIT-Han-Lab (Cited 2024) for SmoothQuant and QServe. MixLLM uses 128 samples with sequence length of 2048. SmoothQuant and QServe uses 64 samples with sequence length of 1024 to prevent OOM.
Metrics. We compare the perplexity (ppl) between all the baselines on wikitext2 and C4 Raffel et al. (2020) datasets. We also compare a set of popular downstream tasks on Llama 3.1 8B, Qwen2.5 7B, and Mixtral 7B v0.3 through lm-eval Gao et al. (2024), including BBH Suzgun et al. (2023), GPQA Rein et al. (2023), MMLU-Pro Wang et al. (2024), MuSR Sprague et al. (2024), ARC challenge Boratko et al. (2018), and HellaSwag Zellers et al. (2019). We conduct the system experiments on NVIDIA A100 (80G) GPUs with CUDA 12.1, except for that Sec.4.3.2 discusses the performance on NVIDIA H100 GPU. We use PyTorch 2.4.1 and transformers 4.45.2.
4.2 Perplexity Evaluation
Tab.1 shows the perplexity on Wikitext2 and C4 dataset for the commonly used open source LLMs, of different baselines. It shows that: 1) Using 4.4 bits of weights with MixLLM can achieve the similar accuracy with 5 bits RTN weight-only quantization, even with 8-bit activation quantization enabled in MixLLM. This is mainly because MixLLM assigns the high-salience output channels with larger bit-widths than the uniform 5-bit solution. 2) As for the weight-activation quantization baselines, MixLLM W4.4A8 shows a comparable accuracy with SmoothQuant with much smaller bit-width (60% of that in SmoothQuant). MixLLM W4.4A8 shows better accuracy than QuaRot and QServe with only 10% larger bit-width. It shows MixLLM achieves a good balance of memory consumption and accuracy. 3) Note that MixLLM W8A8 quantization shows nearly lossless performance compared to the float16 baseline. Besides, the MixLLM W4A8 also outperforms the SOTA QuaRot and QServe for many cases, due to using group-wise quantization for the activation in MixLLM rather than the per-token method in QuaRot and QServe. This is part of the motivation that MixLLM uses group-wise quantization for the activation.
Figure 5: The speedup of two types of single linear layers over torch FP16 baseline on A100 GPU.